OpenStudy (anonymous):

Evaluate t plus 3 when t equals twelve. 4 9 fifteen thirty-six

OpenStudy (anonymous):

we have : t+3=? t=12 (12)+3=?

OpenStudy (anonymous):

ok i am still not getting it

OpenStudy (cggurumanjunath):

WHAT IS t plus 3 @kingdc23 ?

OpenStudy (cggurumanjunath):

CAN U WRITE IT MATHEMATICALLY

OpenStudy (anonymous):

fifteen i was not thinking

OpenStudy (cggurumanjunath):

t plus 3

OpenStudy (cggurumanjunath):

t+3 12+3